# A Cycle Ride to HDR: Semantics Aware Self-Supervised Framework for Unpaired LDR-to-HDR Image Translation

### v1

@@ 5.1 HDR Reconstruction
Ours 31.71 0.941 0.092
Table 2 summarizes the results of evaluation on HDRTV [13], NTIRE [64], HDR-Synth & HDR-Real [51] paired datasets, the most widely used datasets by the state-of-the-art (cf., Supplementary for an extensive qualitative evaluation). In the top part of the table, we compare our method with $11$ supervised learning approaches. The proposed method outperforms those approaches in terms of SSIM, LPIPS, and HDR-VDP-2. In terms of PSNR, it is second best after Ghost-free HDR. In the bottom part of the table, we compare our method with $4$ unsupervised/weakly-supervised/self-supervised learning methods. The proposed method outperforms those approaches on all evaluation metrics (cf., Supplementary for visual depiction of the quantitative HDR reconstruction results). It is worth noting that we used paired datasets in order to compare our method with the state-of-the-art, which predominantly requires paired samples for training. However, in all experiments, our method is trained in a self-supervised fashion using only unpaired samples (cf., Section 3).
To further demonstrate the strength of our approach we extended the train set used in the previous experiment with: 1) The HDR images from DrTMO [19], Kalantari [36], HDR-Eye [43], LDR-HDR Pair [33], and GTA-HDR [7] ($1500$ random samples), i.e., additional $2854$ HDR images; and 2) LDR images from GTA-HDR [7] ($2000$ random samples) and $50\%$ of the LDR images in DrTMO, Kalantari, HDR-Eye, and LDR-HDR Pair, i.e., additional $2677$ LDR images. We kept the test set identical to the previous experiment in order to compare with the state-of-the-art paired data based methods. Given that the only other method that supports this type of unpaired train data is UPHDR-GAN, we perform a direct comparison with that approach in Table 3. The results demonstrate that our method outperforms UPHDR-GAN and sets new state-of-the-art on all metrics (cf., Supplementary for more unpaired data experiments).
@@
Ours 35.11 0.969 0.029 67.88
Ours (mod) 43.01 0.988 0.023 68.21
@@
Table 6: Architecture ablation results with the HDRTV [13], NTIRE [64] and HDR-Synth & HDR-Real [51] datasets. Uo: The U-Net generator used in SingleHDR(W) [45], Uf: The proposed U-Net generator, E: CLIP embedding encoder.
@@
Uf+E+$\mathcal{L}_{\text{con}}$ 36.45 0.911 0.047 66.56
Uf+E+$\mathcal{L}_{\text{sem}}$ 34.21 0.897 0.094 66.12
Uo+E+$\mathcal{L}_{\text{con}}$+$\mathcal{L}_{\text{sem}}$ 35.15 0.953 0.042 67.18
Ours 40.01 0.979 0.021 68.41
5.2 LDR Reconstruction
For a complete evaluation, we also report results for LDR image reconstruction in Table 4 and compare with the state-of-the-art tone-mapping operators $\mu$-law [34], Reinhard’s [68], and Photomatix [28]. The test data is identical to the one used in the HDR reconstruction experiments. The results demonstrate that our method outperforms the other approaches in terms of SSIM and LPIPS. In terms of PSNR, it is second best after Photomatix [68] (cf., Supplementary for more LDR reconstruction experiments).
5.3 Ablation Results
Generators. We propose a novel generator that modifies the U-Net architecture by introducing a feedback mechanism (cf., Section 3.1). The feedback helps in avoiding artifacts in image-to-image translation tasks where huge amounts of training data might make the process of hallucinating details difficult. To test this hypothesis, we first replaced the original U-Net used in the SingleHDR(W) [45] method with our U-Net generator. We see that leads to improvement in all metrics as shown in Table 5. The table also demonstrates the improvement in our model when we use the feedback U-Net instead of the original U-Net of SingleHDR(W) (cf., Supplementary for qualitative results in support of this quantitative evaluation).
@@ 5.3 Ablation Results
Architecture. Our method uses the cycle consistency concept of CycleGAN and introduces many new components to address the LDR $\leftrightarrow$ HDR translation task. The most significant components are the CLIP embedding encoder and feedback U-Net generators (cf., Section 3.1). We perform a thorough ablation on each of the components to highlight the contributions. Table 6 summarizes the results where we add each component to the architecture as we go down. The first row lists the results from the original CycleGAN method. The second row is the CycleGAN utilizing the U-Net generator from SingleHDR(W) [45]. The third row is the CycleGAN utilizing our feedback U-Net generator. This results in a significant improvement in SSIM, LPIPS, and HDR-VDP-2. The fourth row lists the results when we add the CLIP embedding encoder to the architecture. This results in an improvement in all metrics. The fifth and sixth rows depict the improvements after adding contrastive and semantic segmentation objectives in the architecture (cf., Section 3.2). The seventh row depicts the results with all components in our architecture but with the original U-Net from SingleHDR(W). Finally, the last row is the full architecture with all components which achieves the best performance on all metrics.
@@
Loss Functions. We also study the loss functions and their contribution towards the overall results keeping the architectural components constant. Table 7 summarizes the results. The first row is the results from the standard adversarial and cycle consistency loss. In the second row we add the identity loss which improves all metrics (cf., Supplementary for qualitative results in support of $\mathcal{L}_{\text{id}}$). In the third and forth rows we add the contrastive and semantic segmentation loss. This results in an improvement in all metrics (cf., Supplementary for qualitative results in support of $\mathcal{L}_{\text{con}}$ and $\mathcal{L}_{\text{sem}}$). Finally, adding all loss functions achieves the best results on all the metrics.
@@
Table 7: Loss functions ablation results with the HDRTV [13], NTIRE [64] and HDR-Synth & HDR-Real [51] datasets.
@@
We also perform separate tests on contrastive and semantic segmentation loss to examine the contribution of histogram-equalized LDR compared to original LDR images. We summarize these results in Tables 8 and 9. In both cases the loss calculated between the histogram-equalized versions of the LDR (instead of original the LDR) and the recnstructed HDR is better. This is due to the fact that histogram-equalized LDR can reveal semantic information in extremely over/underexposed areas of the original LDR images which in turn improves the extracted CLIP embeddings for $\mathcal{L}_{\text{con}}$ and segmentation maps for $\mathcal{L}_{\text{sem}}$.
Table 8: Contrastive loss ablation results with the HDRTV [13], NTIRE [64] and HDR-Synth & HDR-Real [51] datasets.
@@
LDRHist 40.01 0.979 0.021 68.41
Table 9: Segmentation loss ablation results with the HDRTV [13], NTIRE [64] and HDR-Synth & HDR-Real [51] datasets.
[…]
LDR 38.89 0.956 0.033 66.92
LDRHist 40.01 0.979 0.021 68.41
6 Conclusions
Most current state-of-the-art methods for HDR reconstruction (inverse tone-mapping) require high-quality paired {LDR,HDR} datasets for model training. Consequently, the quality of the state-of-the-art in inverse tone-mapping depends on the quality of the available paired {LDR,HDR} datasets. Additionally, only a few approaches utilize semantic and contextual knowledge guidance in the reconstruction process which could significantly improve the quality of output HDR images.
[…]

### v2

@@ 5.2 LDR Reconstruction
GlowGAN-prior [94] 32.03 0.936 0.108
LLM-HDR (Ours) 31.81 0.942 0.091
5.3 Ablation Results
Generators. We propose a novel generator that modifies the U-Net architecture by introducing an artifact-aware feedback mechanism (Sec. 3.1). The feedback helps in avoiding artifacts in image-to-image translation tasks where huge amounts of training data might make the process of hallucinating details difficult. To test this hypothesis, we first replaced the original U-Net used in the SingleHDR(W) [55] method with our U-Net generator. We see that this leads to improvements in all but one metrics as shown in Tab. 6. The table also demonstrates the improvement in our model when we use the feedback U-Net instead of the original U-Net of SingleHDR(W). Fig. 3 illustrates the improvement in SingleHDR(W) [55] when we use the proposed feedback U-Net instead of the original U-Net of SingleHDR(W). The original U-Net produces many artifacts in the output HDR images whereas our modified version with feedback reconstructs artifact-free HDR images. Sec. S3.3 provides extensive ablation experiments for the proposed architecture and loss functions.
Figure 3: Comparison of the SingleHDR(W) [55] U-Net with and without our feedback mechanism on images from the DrTMO [23].
6 Conclusion
[…]
@@ 6 Conclusion
LLM-HDR 35.11 0.969 0.029 67.88
LLM-HDR (mod) 43.01 0.988 0.023 68.21
Supplementary Material
S1 Related Work
This section complements Sec. 2 in the main paper and includes an overview of non-learning approaches for HDR reconstruction. The most popular approach in this category consists of multi-exposed LDR fusion to achieve high dynamic range in the output image [42]. This approach involves image feature alignment, calculating weights for feature mapping on the basis of image characteristics, and fusing the images on the basis of the weights to get the most appropriate exposures for each part of the image. Another approach is using histogram equalization [20] that allows for contrast and brightness levels re-adjustment in the over/underexposed regions of the image. Gradient domain manipulation [52] techniques enhance the granular details of an LDR image and expand its dynamic range. Pyramid-based image fusion [46] is a technique where image features are extracted into Laplacian pyramids [46] and then fused for the HDR. Some methods use Retinex-based [48, 71] tuning to effectively produce and display HDR images on consumer screens [48, 71]. Intensity mapping function-based approaches [110, 21, 80] use transformations on image pixels either by using Logarithmic mapping or Gamma correction to stretch the intensity range or enhance lower intensity pixels. The methods in this category, although time-efficient, face some issues including ghosting effects, halo artifacts, and blurring in the reconstruction, and they do not generalize well for variety of input LDR images.
S2 Method
This section complements Sec. 3 in the main paper. It provides a concise background on the key concepts employed in our method, i.e., cycle consistency [117], Large Language Models [75, 105], contrastive language-image pretraining [81, 62, 112], and image semantic segmentation [50]. The section also offers implementation details for the architecture modules.
S2.1 Background
Cycle Consistency. The concept of cycle consistency in image-to-image [117] and video-to-video [17] translation tasks is a powerful and elegant design for unpaired data. Previously, cycle consistency has been applied to paired LDR $\leftrightarrow$ HDR translation [106], where the authors used $3$ LDR images with different exposures as input to reconstruct an HDR. The reconstructed HDR is then used by $3$ different generators to reconstruct the original $3$ multi-exposed LDR images. However, the focus of our work is on the more general task of unpaired LDR $\leftrightarrow$ HDR translation. Given the images $\{x_{i}\}_{i=1}^{N}$ where $x_{i}\in X$ and $\{y_{j}\}_{j=1}^{M}$ where $y_{j}\in Y$, the goal is to design two generators, i.e., $G_{Y}$ and $G_{X}$ such that, $G_{Y}(x)$rightarrow $y$ and $G_{X}(y)$rightarrow $x$. Fig. S1 depicts this concept for the LDR-to-HDR translation task.
Figure S1: $G_{Y}$ translates LDR to HDR images and $G_{X}$ HDR to LDR. The LDR images are from the LDR-HDR pair [40] dataset.
LLM and VLM. Large Language Models [75] and Vision-Language Models [114, 105] are trained on vast amounts of textual and/or visual data (i.e., images and videos) and achieve excellent zero-shot performance. We leverage the text and visual perception capabilities of an LLM to evaluate the reconstruction of our method and refine it with better awareness of artifacts and over/underexposed regions.
Contrastive Language-Image Pretraining. CLIP [81] is a method that finds the closeness between visual and textual data in the embedding space. The method is trained on image-text pairs to associate images with their descriptions. The CLIP loss maximizes the cosine similarity between correct (positive) pairs and minimize the similarity between incorrect (negative) pairs. We extract the CLIP embeddings from both the LDR and reconstructed HDR image pairs and maximize their cosine similarity in the embedding space for preservation of semantic information across domains. We also combine the CLIP embeddings extracted from the LDR and reconstructed HDR to guide the decoder of the generators in the next step of the reconstruction.
Semantic Segmentation. Semantic segmentation [50, 83, 59] is a vision problem where each pixel of an image or video frame is segmented into meaningful sub-parts such as objects, roads, water bodies, buildings, humans, and animals. Segment Anything (SAM) [50] is a state-of-the-art method for semantic segmentation of images. Segmented objects in a scene not only inform the number and location of semantically meaningful concepts, but also provide a holistic semantic description of the scene. Therefore, we can use that information to improve image reconstruction. We introduce a loss on the basis of the Mean Intersection over Union metric to preserve the image integrity (based on the segmented objects) between the two domain $X$ and $Y$.
Figure S2: Overview of the proposed generators based on our novel feedback based U-Net architecture. Left part (blue) is the encoder and right part (red) is the decoder. The decoder is inside the feedback iteration loop.
Figure S3: Overview of the proposed LLM module. Q&A sessions include: User prompt: “Tell me if there are any synthesis artifacts in the given scene or not. Must response with 1) Yes or No only, 2) If Yes, can you get me the saliency maps of the artifacts (in red), overexposed (in blue) and underexposed (in yellow) areas of this image?”, System: “Outputs the saliency maps for the detected artifacts in red, overexposed pixels in blue, and underexposed pixels in yellow.”
S2.2 Architecture Modules
Generators. The U-Net of each generator consists of $7$ levels. Each level consists of $2$ Conv layers with $3\times 3$ kernel followed by a ReLU activation and a BN layer. The input images of size $512\times 512$ are first converted into $32$-dimensional features. The features are doubled in each level of the encoder until reaching size of $512$. The decoder levels then up-sample the data by performing channel-wise concatenation with the extracted feature from previous levels of the encoder. The intuition behind using feedback mechanism is to refine the decoder layers of the generator based on the input from encoder layers and the feedback mechanism itself. The output of each iteration of the feedback mechanism is not only fed into the first level of the decoder block, but also sent back through the hidden state to guide its next input. Feedback mechanisms perform better than simple feed-forward networks with less trainable parameters [63]. The feedback network consists of $3$ dilated dense blocks each with dilation rate of $3$. The dilated blocks are made of $3\times 3$ Conv layers followed by ReLU activation function. Each dilated block contains two $1\times 1$ feature compression layers at the start and end. Dilation enhances the network’s receptive field which helps to capture broader context from the input features without increasing the computational complexity. Finally, the feedback starts with a $1\times 1$ Conv and ends with $3\times 3$ Conv for compression and final processing, respectively. Fig. S2 depicts and overview of the proposed generators architecture.
Figure S4: Depiction of the LLM-based loss $\mathcal{L}_{\text{llm}}$. Q&A sessions include: User prompt: “Tell me if there are any synthesis artifacts in the given scene or not. Must response with 1) Yes or No only, 2) If Yes, return the [artifact areas number of pixels, total number of pixels]. Also highlight the artifact areas with pink. Similarly find if there are any over/under exposure areas in the scene and return the [pixels in overexposure, pixels in underexposure areas] and highlight the overexposed areas in blue and underexposed areas in green.”, System: “Outputs the number of pixels with the pixel list (in .txt files) for the detected artifacts in pink, overexposed pixels in blue, and underexposed pixels in green.”
Figure S5: Depiction of the contrastive loss $\mathcal{L}_{\text{con}}$. Positive (green) and negative (red) pairs in a batch. We use a histogram-equalized version of the LDR processed using the OpenCV function equalizeHist [10].
Figure S6: Depiction of the semantic segmentation loss $\mathcal{L}_{\text{sem}}$. We use Segment Anything (SAM) [50] to generate segmentation classes in the histogram-equalized LDR and reconstructed tone-mapped HDR images.
Figure S7: Depiction of the cycle consistency loss $\mathcal{L}_{\text{cyc}}$ using an image from the DrTMO [23] dataset.
Discriminators. The discriminators consist of $5$ Conv layers. We input the reconstructed LDR/HDR along with a real (unpaired) LDR/HDR in the respective discriminator. The Conv layers have $4\times 4$ filters and stride of $2$ for the first $3$ layers and $1$ henceforth. The output of the first layer has features of size $64$, which is doubled in each of the next layers until $512$. The output of the fifth layer is of size $1$. The first four layers have LeakyReLU activation with the final layer using a Sigmoid activation for outputting the probabilities of a particular image being real or fake. The training process is stabilized using BN following all the Conv layers except the first one.
LLM. The artifact-aware saliency map $LLM_{\text{af}}$ is extracted as feature of size $512$ with $1\times 1$ Conv and we perform an element-wise multiplication with the output features of the bottleneck layer (i.e., layer $4$) of the U-Net generators $G_{Y}$ and $G_{X}$ which goes into the feedback mechanism. The exposure-aware saliency map $LLM_{\text{ox}}$ (for overexposure) (which goes into the generator $G_{Y}$ only) is extracted as feature of size $256$ with $1\times 1$ Conv and fused with the output features of layer $3$ henceforth concatenated with the encoder embeddings which goes into the bottleneck layer. Finally, $LLM_{\text{ux}}$ (for underexposure) is extracted as feature of sizes $64$, $128$, and $256$, with $1\times 1$ Conv and fused with the skip connections in each level of the U-Net generator $G_{Y}$ (i.e., level 1, 2, and 3). Fig. S3 depicts the overview of the proposed $LLM$ module.
Encoders. The encoders extract embeddings of size $512$ from the input and output domain images. Given the input size of the generators, we re-project these embeddings to a size of $256$ using $1\times 1$ Conv. Then we add the embeddings and concatenate them with the input features of the bottleneck layer (i.e., layer $4$) of our feedback-based U-Net generators.
S2.3 Loss Functions
Figs. S4, S5, S6 and S7 illustrate the formulation of the LLM-based loss, contrastive, semantic segmentation, and cycle consistency functions, respectively.
S2.4 Implementation Details
The proposed method is implemented in PyTorch [78] on Ubuntu 20.04.6 LTS workstation with Nvidia Quadro P5000 GPU with 16GB memory, Intel® Xeon® W-2145 CPU at 3.70GHz with 16 CPU cores, 64GB RAM, and 2.5TB SSD. We used Adam optimizer [49] to train the models for $170$ epochs. Batch size of $1$ is expected to work well in cycle consistency models [117] however, since we are using contrastive loss on positive and negative pairs, the batch size is set to $4$. We used learning rate of $4\times 10^{-4}$ for the generators and $2\times 10^{-4}$ for the discriminators. The learning rate is kept constant for the first $100$ epochs and subsequently set to decay linearly to 00. For all models, we resized the input images to $512\times 512$. The HDR images used in the text are tone-mapped with Reinhard’s operator [84].
S3 Results
This section complements Sec. 5 in the main paper. It provides additional quantitative evaluation of our method for both HDR and LDR reconstruction. The section also includes an extensive qualitative and ablation results.
S3.1 HDR Reconstruction
Tab. S1 provides the results from an additional mixed data experiment. The train set used in this experiment is: 1) The HDR images from DrTMO [23], Kalantari [44], HDR-Eye [51], LDR-HDR Pair [40], and GTA-HDR [7] ($1500$ random samples), i.e., $2854$ HDR images; and 2) LDR images from GTA-HDR [7] ($2000$ random samples) and $50\%$ of the LDR images in DrTMO, Kalantari, HDR-Eye, and LDR-HDR Pair, i.e., $2677$ LDR images. In this dataset, a total of $1177$LDR,HDR} pairs overlap. Given that the only other methods that support this type of unpaired train data are UPHDR-GAN [61] and GlowGAN-ITM [94], we perform a direct comparison with these two approaches. The results demonstrate that our method again outperforms UPHDR-GAN and GlowGAN-ITM on all metrics.
Table S1: HDR reconstruction results. Comparison with unsupervised learning methods trained on mixed datasets.
[…]
UPHDR-GAN [61] 41.98 0.975 0.046 65.54
GlowGAN-ITM [94] 32.18 0.905 0.066 61.89
LLM-HDR (Ours) 42.01 0.988 0.023 68.81
Figs. S12, S13, S14, S15, S16, S17 and S18 illustrate the quality of the HDR images reconstructed with our method. Our method mitigates artifacts such that the resulting HDR closely resembles the ground truth HDR. The finer textures, color hues, and shades are well preserved compared to the other methods. The granular level details in the bright and dark regions are also well reconstructed.
S3.2 LDR Reconstruction
Tab. S2 provides the results from an additional LDR reconstruction experiment. In this case, our method is trained with the data specified in the previous section. Similar to the LDR reconstruction experiment in the main paper, we compare the model with the state-of-the-art tone-mapping operators $\mu$-law [43], Reinhard’s [84], Photomatix [34], and GlowGAN-prior [94]. The results again demonstrate that our method outperforms the other approaches in terms of SSIM and LPIPS. In terms of PSNR, it is third best after Photomatix and GlowGAN-prior.
Table S2: LDR reconstruction results. Comparison with the state-of-the-art tone-mapping operators and GlowGAN-prior [94].
[…]
$\mu$-law operator [43] 28.68 0.891 0.311
Reinhard’s operator [84] 30.12 0.903 0.342
Photomatix [34] 32.11 0.922 0.136
GlowGAN-prior [94] 31.91 0.931 0.112
LLM-HDR (Ours) 31.11 0.939 0.095
@@
Table S3: Architecture ablation results with the HDRTV [16], NTIRE [79] and HDR-Synth & HDR-Real [64] datasets. Uo: The U-Net generator used in SingleHDR(W) [55], Uf: The proposed U-Net generator, E: CLIP embedding encoder, LLM: LLM module for artifact-aware and exposure-aware maps.
@@
Uf+E+LLM 30.15 0.887 0.101 64.41
Uf+E+LLM+$\mathcal{L}_{\text{con}}$ 36.45 0.911 0.047 66.56
Uf+E+LLM+$\mathcal{L}_{\text{sem}}$ 34.21 0.897 0.094 66.12
Uf+E+LLM+$\mathcal{L}_{\text{llm}}$ 36.01 0.907 0.046 66.72
Uf+E+LLM+$\mathcal{L}_{\text{con}}$+$\mathcal{L}_{\text{sem}}$ 40.01 0.979 0.021 68.41
Uo+E+LLM+$\mathcal{L}_{\text{con}}$+$\mathcal{L}_{\text{sem}}$+$\mathcal{L}_{\text{llm}}$ 35.15 0.953 0.042 67.18
Uf+E+LLM+$\mathcal{L}_{\text{con}}$+$\mathcal{L}_{\text{sem}}$+$\mathcal{L}_{\text{llm}}$ 40.11 0.979 0.020 68.51
S3.3 Ablation Results
@@
Architecture. Our method uses the cycle consistency concept of CycleGAN and introduces many new components to address the LDR $\leftrightarrow$ HDR translation task. The most significant components are the LLM module, artifact-aware feedback and exposure-aware skip connection-based U-Net generators, and the CLIP embedding encoders (Sec. 3.1). We perform a thorough ablation on each of the components to highlight the contributions. Tab. S3 summarizes the results where we add each component to the architecture as we go down. The first row lists the results from the original CycleGAN method. The second row is the CycleGAN utilizing the U-Net generator from SingleHDR(W). The third row is the CycleGAN utilizing our feedback U-Net generator. This results in a significant improvement in SSIM, LPIPS, and HDR-VDP-2 (Q-Score). The fourth row shows the results when we add the LLM module to the architecture which shows some improvement in the PSNR, LPIPS, and Q-Score. The fifth row lists the results when we remove the LLM module and add the CLIP embedding encoders to the architecture. This results in an improvement in PSNR, SSIM, and Q-Score. The sixth row depicts good improvement in almost all the metrics as we add both CLIP encoders and LLM module to the architecture. The seventh, eight, and ninth rows depict the improvements after adding contrastive, semantic segmentation, and LLM-based objectives (Sec. 3.2) in the architecture, respectively. The tenth row shows the improvements when we add only the contrastive and semantic segmentation objectives in the architecture. The eleventh row depicts the results with all components in our architecture but with the original U-Net from SingleHDR(W). The last row is the full architecture with all components which achieves the best performance on all metrics.
@@
Loss Functions. We also study the loss functions and their contribution towards the overall results keeping the architectural components constant. Tab. S4 summarizes the results. The first row is the results from the standard adversarial and cycle consistency loss. In the second row we add the identity loss which improves all metrics. In the third, fourth, and fifth rows we add the contrastive, semantic segmentation, and LLM-based objectives. This results in an improvement in all metrics. The sixth row shows the results when we add contrastive and semantic losses together which gives significant improvement in the metrics. Finally, adding all loss functions (seventh row) achieves the best results on all metrics.
Fig. S9 visually validates the contribution of the identity loss $\mathcal{L}_{\text{id}}$ illustrating that it helps in recovering the hue and shades of original image in the reconstruction. Figs. S10 and S11 demonstrate similar results for the contrastive and semantic segmentation objectives. The contrastive loss $\mathcal{L}_{\text{con}}$ can recover high-level attributes such as color and texture information more accurately compared to our method without this loss component. We also observe that the semantic segmentation loss $\mathcal{L}_{\text{sem}}$ recovers the low-level attributes such as object boundaries and edges more vividly than our method without this loss component. These high/low-level attributes are of vital importance to downstream applications like object recognition.
@@
Table S4: Loss functions ablation results with the HDRTV [16], NTIRE [79] and HDR-Synth & HDR-Real [64] datasets.
@@
+$\mathcal{L}_{\text{id}}$+$\mathcal{L}_{\text{con}}$+$\mathcal{L}_{\text{sem}}$+$\mathcal{L}_{\text{llm}}$ 40.11 0.979 0.020 68.51
We also perform separate tests on contrastive and semantic segmentation loss to examine the contribution of histogram-equalized LDR compared to original LDR images. We summarize these results in Tabs. S6 and S5. In both cases the loss calculated between the histogram-equalized versions of the LDR (instead of original the LDR) and the recnstructed HDR is better. This is due to the fact that histogram-equalized LDR can reveal semantic information in extremely over/underexposed areas of the original LDR images which in turn improves the extracted CLIP embeddings for $\mathcal{L}_{\text{con}}$ and segmentation maps for $\mathcal{L}_{\text{sem}}$.
Table S5: Semantic segmentation loss ablation results with the HDRTV [16], NTIRE [79] and HDR-Synth & HDR-Real [64] datasets.
[…]
LDR 38.89 0.956 0.033 66.92
LDRHist 40.11 0.979 0.020 68.51
Table S6: Contrastive loss ablation results with the HDRTV [16], NTIRE [79] and HDR-Synth & HDR-Real [64] datasets.
@@
LDRHist 40.11 0.979 0.020 68.51
Finally, we experiment with the weights for the loss components with respect to PSNR. We perform tests on the temperature parameter $\tau$ of the contrastive loss and the weight parameters $\alpha$ and $\beta$ for the loss components $\mathcal{L}_{\text{con}}$ and $\mathcal{L}_{\text{sem}}$. Fig. S8 shows the results. The temperature parameter $\tau$ gives the optimal PSNR with $\tau=0.08$ while the result degrades if we increase the value further. If we decrease the value, it gives a sub-optimal PSNR at $\tau=0.07$, so we fix the value to $\tau=0.08$ in all experiments. The values of $\alpha$ and $\beta$ give optimal PSNR for $\alpha=2$ and $\beta=2$. The weights for the LLM-based loss are set to $3$, $2$, and $1.5$ for the artifact, overexposed, and underexposed components, respectively. We conducted ablation experiments on these weights considering values from $1$ to $5$ and found that the selected combination of weights yield the best PSNR.
Figure S8: Ablation results for the temperature parameter $\tau$ (a), and loss component weights $\alpha$ (b) and $\beta$ (c).
S4 Limitations and Future Work
Our method is designed for LDR $\leftrightarrow$ HDR bi-directional translation for images and can be further extended to video-based LDR $\leftrightarrow$ HDR translation with some modifications in the generators (e.g., using special ViT [32] or Timesformer [9]) models. Diffusion-based cycle consistency approach can also be explored for video-based LDR/HDR translation. Further research can also explore multi-task architectures along with dynamic range translation such as deblurring, denoising, super-resolution, and demosaicing tasks. The pre-processing step in our loss calculations for contrastive and semantic segmentation loss employs histogram equalization, which can be further investigated with methods like Gamma correction [58] or histogram matching [18]. Another direction can be the integration of no-reference quality assessment [4] into the framework which will enable parallel training and prediction of the reconstructed LDR/HDR along with its quality score. More advanced LLM-based formulation such as explainability of the generation process and object’s material based specular reflections can also be introduced in the future. Moreover, the Differentiable Ray Tracing technology [24] used to generate realistic HDR images/videos in video games [7] to simulate the physics behind light, specular reflections and shadows can also be explored using deep learning concepts like Nvdiffrast [53] and Neural Radiance Fields (NeRFs) [73] in 2D domain.
Figure S9: Ablation results for $\mathcal{L}_{\text{id}}$ with images from the HDR-Synth & HDR-Real dataset [64] and LDR-HDR pair datasets [40].
Figure S10: Ablation results for $\mathcal{L}_{\text{con}}$ with images from the LDR-HDR pair [40] dataset.
Figure S11: Ablation results for $\mathcal{L}_{\text{sem}}$ with images from the HDR-Synth & HDR-Real dataset [64] dataset.
Figure S12: Examples of HDR images reconstructed with our method and recent state-of-the-art.
Figure S13: Examples of HDR images reconstructed with our method and recent state-of-the-art.
Figure S14: Examples of HDR images reconstructed with our method and recent state-of-the-art.
Figure S15: Examples of HDR images reconstructed with our method and recent state-of-the-art.
Figure S16: Examples of HDR images reconstructed with our method and recent state-of-the-art.
Figure S17: Examples of HDR images reconstructed with our method and recent state-of-the-art.
Figure S18: Examples of HDR images reconstructed with our method and recent state-of-the-art.